container A contains 3 times as much water as container b. if you pour 1.25 quarts from a to b. now container a has only twice as much water than b.
what were the amounts in a and b to start??

Answer by lwsshak3(11628)

container A contains 3 times as much water as container b. if you pour 1.25 quarts from a to b. now container a has only twice as much water than b.
what were the amounts in a and b to start?
**
let x=quarts of water in container b to start
3x =quarts of water in container a to start
..
3x-1.25=2(x+1.25)
3x-1.25=2x+2.5
x=2.5+.125=3.75
3x=3*3.75=11.25
ans:
quarts of water in container b to start=3.75 quarts
quarts of water in container a to start=11.25 quarts